OpenStudy (anonymous):

A neighborhood has provided trash cans to residents to dispose of their waste. All trash cans have a diameter of 1.5 meters and a height of 1.25 meters. What is the minimum number of trash cans required to dispose of 10 cubic meters of waste? 14 13 5 4

OpenStudy (ghazi):

volume of each trash = pi*r^2*length = 3.14*(1.5/2)^2*1.25=2.207m cube

OpenStudy (ghazi):

volume of one trash can= 2.207 m^3 so for 1m^3 you need= 1/2.707 cans therefore for 10 you need number of cans = 1/2.707*10=3.69 cans which is 4 cans

OpenStudy (ghazi):

@devibabyy66 is it clear?

OpenStudy (anonymous):

yes its clear now thank you so much!

OpenStudy (ghazi):

:)